The reciprocal of two plus the reciprocal of four equals the reciprocal of 2x. Find x.
Answer by checkley71(8403)

1/2+1/4=1/2X
(2+1)/4=1/2X
3/4=1/2X NOW CROSS MULTIPLY.
3*2X=4
6X=4 NOW DIVIDE BOTH SIDES BY 6.
X=4/6 NOW REDUCE.
X=2/3 ANSWER.